OpenStudy (anonymous):

how do i convert 112 degree to radian

OpenStudy (anonymous):

use (pi)^c = 180 degree

OpenStudy (anonymous):

are you there? i might be able to help

OpenStudy (anonymous):

basically i figured this to be really easy when you use this formula

OpenStudy (anonymous):

i did figure it out thank you

OpenStudy (anonymous):

oh right when im about to draw the equation xD

OpenStudy (anonymous):

oh im sorry but i do have a nother question

OpenStudy (wolf1728):

There are 2*PI radians in every circle so 2*PI = 360°. Therefore, 360° / 2*PI equals the number of degrees in a radian. 360° / 2*PI = 360° / 6.28318530717959 = 57.2957795130823 degrees in every radian.

OpenStudy (wolf1728):

So, 112° = 112 / 57.296 = 1.9548 radians

OpenStudy (anonymous):

how do i find the measure of a complementary angle to pi/18 radian

OpenStudy (anonymous):

okay first of all, do you know what a complementary angle is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i know how to measure supplemenraty angle, but im confused with complementary

OpenStudy (anonymous):

okay supplementary angles are added up to 180 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Complementary angles should be added upto 90

OpenStudy (anonymous):

oh, how would i write that out

OpenStudy (anonymous):

so what you want to do is this.

OpenStudy (anonymous):



OpenStudy (anonymous):

thats the problem right?

OpenStudy (anonymous):

you will convert radians to degrees

OpenStudy (anonymous):

yes

OpenStudy (anonymous):